OpenStudy (anonymous):

find the surface area of the sphere with the given dimension . leave answers in term of pi. Diameter: 16cm

OpenStudy (anonymous):

\[256cm \pi\]

OpenStudy (anonymous):

\[128cm \pi\]

OpenStudy (anonymous):

\[32cm \pi\]

OpenStudy (anonymous):

\[1024cm \pi\]

OpenStudy (anonymous):

@antwone

OpenStudy (mathstudent55):

Do you know the formula for the surface area of a sphere?

OpenStudy (mathstudent55):

Surface area of sphere: \(A = 4 \pi r^2\) d = 16 cm. The radius is half the diameter. Now replace r with 8 cm, and calculate A.

OpenStudy (anonymous):

@mathstudent55 :i got 100.48 ...

OpenStudy (mathstudent55):

Don't use 3.14 for pi. Leave pi as pi. If you use the formula, you get: \(A = 4 \times \pi \times (8~cm)^2\) First, calculate \(8^2\). Then multiply that by 4. The answer is the number you get times pi.